# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ I-D5 Activities in 2022
Many different types of activities occurred in January. The project TiC6G was launched by the University of Glasgow, with a group of academic and industrial partners, aiming to test prototype devices at THz frequencies using the university’s new cutting-edge laboratory equipment. The Japanese operator Nippon Telegraph and Telephone (NTT) reported that it will build a 6G pilot network for the Osaka World Expo 2025. The Chinese Purple Mountain Laboratories, together with Fudan University and China Mobile, reported that it had obtained a data rate of 206.25 Gbit/s for a wireless THz transmission in a laboratory environment. The Next G Alliance published a white paper titled ”Green G: The Path toward Sustainable 6G”. The University of Oulu announced a partnership with Jio Estonia to study 6G technologies, particularly for non-terrestrial communications, holographic radio, and three-dimensional (3D) connected intelligence.
@@
In February, Keysight Technologies announced a partnership with Samsung to develop, test, and verify 6G technologies. The Next-Generation Mobile Networks (NGMN) Alliance published a white paper on 6G use cases, categorizing them into four main classes, i.e., enhanced human communications, enhanced machine communications, enabling services, and network evolution. At the Mobile World Congress (MWC) 2022, LTE showcased its cutting-edge RIS technology, including four different prototypes, i.e., liquid crystal RIS, PIN diode RIS, transparent planar RIS, and transparent flexible RIS. At the same venue, the US technology company VMware reported its collaboration with German research institutes to accelerate 6G research, focusing on the fusion of cloud, networking, and AI technologies.
@@
In April, Kyocera, a Japanese technology firm, announced that it has developed a transmissive metasurface that can be used to extend coverage in 5G and 6G networks by redirecting radio signals to avoid obstacles. In May, an MoU was signed between the European 6G Smart Networks and Services Industry Association (6G-IA) and the Japanese Beyond 5G Promotion Consortium to accelerate the development of 6G networks and deepen the cooperation between Europe and Japan. Samsung released a new white paper, ”6G Spectrum: Expanding the Frontier”, discussing spectrum policies and potential frequency bands for 6G. The Finnish leading academic and industrial organizations established a national 6G coalition, called 6G Finland, aiming to strengthen national 6G cooperation, form new international partnerships, and promote Finland’s 6G expertise on a global scale. LG demonstrated its 6G relevant full-duplex and power amplifier technologies at the IEEE ICC 2022. The Next G Alliance and the Japanese Beyond 5G Promotion Consortium agreed on 6G cooperation by signing an MoU in May.
@@
In June, the Next G Alliance published a white paper titled ”6G Applications and Use Cases”, focusing on four categories, i.e., networked robotics, XR, distributed communication and sensing, and personalized user experiences. Nokia, NTT, and NTT Docomo formed a 6G partnership to demonstrate an AI/ML-enhanced air interface and sub-THz communication at 140 GHz spectrum. The Singaporean Nanyang Technological University announced the use of Keysight’s test and measurement solutions to verify its on-chip THz electronic-photonic 6G device technologies. The Japanese companies NTT Docomo, Fujitsu, and NEC reported to perform joint 6G indoor and outdoor trials. 6G-IA and the Chinese IMT-2030 Promotion Group signed an MoU for 6G cooperation. The Open RAN (O-RAN) Alliance reported a launch of its Next Generation Research Group, focusing on the utilization of O-RAN technologies in 6G.
@@
In July, LG Uplus and Nokia signed an MoU for collaboration in 5G/6G standardization. Nokia announced its leadership in a German 6G light house project 6G-ANNA, aiming to promote 6G development and standardization in Germany. The Next G Alliance published a white paper on the technologies required to meet the demands of its 6G vision. The UK government reported to grant 25 million pounds funding for the research and development work on 6G technologies. In August, the Next G Alliance and European 6G-IA signed an MoU for the agreement of 6G research collaboration, including joint workshops, webinars, and trials. While already supporting 6G research at the University of Texas at Austin, the Northeastern University, the University of Surrey, and Viavi Solutions announced a new funding program to advance worldwide 6G development in academia and industry. The Next G Alliance reported the establishment of a 6G research council, aiming to promote cooperation between academia and industry and align North America’s 6G strategy with the US and Canadian governments.
In September, LG reported its final THz communication test at the Fraunhofer Heinrich-Hertz Institute in Berlin, with a successful information transfer over a 320 meter outdoor link at a frequency range between 155 and 175 GHz. The Finnish University of Oulu and Japanese National Institute of Communication and Technology (NICT) announced a cooperation in B5G/6G research. Rohde &\&& Schwarz and China Mobile reported their partnership in the development of ISAC solutions for 6G. The Singaporean Infocom Media Development Authority and the Singapore University of Technology and Design revealed their collaboration to establish a 6G Research and Development (R&\&&D) laboratory in Singapore, being the first in kind in Southeast Asia. Nokia informed its 5G/6G collaboration with Vodafone New Zealand.
@@
In October, Keysight Technologies and the Indian Institute of Technology, Madras signed an MoU, expanding their cooperation from 5G to 6G design. Deutsche Telekom reported its leadership in the German government-funded project 6G-TakeOff, aiming to develop an architectural framework for the integration of terrestrial and non-terrestrial networks in 6G. The European Commission granted funding for Hexa-X-II, the second phase of the EU’s 6G flagship program, with Nokia continuing in lead. 35 research projects were chosen in the EU’s 250 million euros call to develop advanced 5G and future 6G technologies, with the goal of strengthening the European expertise in the development of future mobile networks. Ericsson and the University of Texas at Austin deepen their collaboration into 6G-empowered XR. VMware announced that it is launching the Next G-AI Research and Innovation Center in Montreal, Canada, gathering multidisciplinary expertise to advance the technological development of 5G and 6G systems. Samsung Electronics reported the founding of a new 6G research group in the UK, which is part of the Samsung’s global 6G development strategy. The German government-funded project 6G NeXt was kicked-off, with Deutsche Telekom in lead, and the goal of developing a flexible infrastructure platform to study and test future XR applications.
@@
In November, ITU-R WP 5D completed its work on the first report on IMT-2030, i.e., ”IMT-2030 Future Technology Trends” [35]. It was reported that India’s Department of Telecommunications suggests opening up the frequency range from 95 GHz to 3 THz for experimenting and testing of 6G technologies and products. Nokia announced its leadership in the German government-supported KOMSENS-6G project, which focuses on the integration of sensing and communication. The Japanese Beyond 5G Promotion Group and Northeastern University agreed on a 6G collaboration by signing an MoU to foster the development of 6G networks. NTT Docomo and SK Telecom announced their partnership in the fields of smart-life and metaverse, searching for opportunities to jointly produce immersive XR content. Ericsson reported its ten-year 6G research initiative in the UK, focusing on the areas of AI, cognitive networks, and network security. Nokia announced the launch of a new 5G/6G R&\&&D center in Portugal. The UK government reported a 110 million pound grant for 5G/6G research and development. 28 million is dedicated to the collaboration of academia (Universities of York, Bristol, and Surrey) and industry (Nokia, Ericsson, and Samsung), while 80 million is invested in a new 5G/6G R&\&&D laboratory. All types of physical, hybrid, and virtual 6G events were held in 2022. Numerous IEEE special issues on 6G were released. Moreover, many 6G white papers, visions, surveys, and books were written.
@@ I-D6 Activities in 2023
In January, the European Hexa-X-II flagship project (i.e., a continuation of Hexa-X) was launched. Europe’s 6G-IA and the European Telecommunications Standards Institute (ETSI) agreed on a close 5G/6G cooperation with each other by signing an MoU, with the goal of advancing 6G pre-standardization activities. The ATIS and O-RAN Alliance agreed on a partnership via an MoU to foster the development of O-RAN technologies. Rohde &\&& Schwarz announced its channel measurement campaign conducted in urban micro outdoor and indoor scenarios at frequencies of 158 GHz and 300 GHz, contributing to the ITU-R report titled ”Technical feasibility of IMT in bands above 100 GHz”. Finland’s University of Oulu and South Korea’s Electronics and Telecommunications Research Institute announced the launch of their new collaboration project, 6GBRIDGE-6GCORE, aiming to develop a service-centric 6G system architecture.
@@
In February, Ericsson reported its co-leadership with Sweden’s KTH Royal Institute of Technology to coordinate a new 6G research project, called DETERMINISTIC6G, which focuses on developing technological enablers for time-sensitive communications applicable to diverse verticals. The University of Sheffield announced the opening of the new UK Research and Innovation National 6G Radio Systems Facility, with top-notch 6G R&\&&D capabilities and support from more than 40 companies and academic organizations. Nokia, NTT Docomo, and NTT reported two major milestones achieved on the evolution road toward 6G, including sub-THz communication (25 Gbit/s at 144 GHz) and the integration of AI into air interface (ML-based waveform), implemented as proof of concepts at Nokia Bell Labs in Germany. Finland’s 6G Flagship and Brazil’s research institute Inatel announced their partnership in developing advanced 6G solutions for rural and remote areas. At the MWC 2023, Nokia demonstrated network sensing capabilities using its prototype radio equipment. Also at the MWC, Bosch reported its collaboration with Nokia on 6G-based IoT solutions for industry 4.0 applications. Taiwan’s operator Chunghwa Telecom and Ericsson signed an MoU at the MWC, joining forces to advance 5G and 6G standardization.
@@
In March, India’s government unveiled a national 6G initiative, targeting to launch 6G networks in India by 2030. Anritsu and Danish Aalborg University announced their collaborative 6G project on channel sounding for joint communication and sensing at mmWave and sub-THz frequencies. In April, the Spanish operator Telefonica, NEC, Bluespecs, and IMDEA Research Institute reported the launch of the ENABLE-6G project, funded by the EU and Spain’s government. The project aims to develop mechanisms to tackle the challenges 6G will encounter in terms of performance, energy efficiency, network sensing, and privacy/security. Keysight Technologies, the University of Surrey, and the National Physical Laboratory demonstrated over 100 Gbit/s data rates at a 300 GHz carrier frequency using their new sub-THz 6G testbed.
@@
In May, the Next G Alliance released a white paper on the 6G roadmap for vertical industries. The French technology company Capgemini announced the opening of a new 6G research lab in India to enable testing and experimentation of novel 6G solutions. Brazil’s government reported the allocation of 36 million dollars to three research centers to advance the development of 5G, 6G, and O-RAN technologies. Rohde &\&& Schwarz and the French Institute of Electronics, Microelectronics and Nanotechnology announced their cooperation in THz communication research, with successful communication over 645 meters using a THz backhaul link operating at a frequency of 300 GHz.
@@
In June, the Next G Alliance published a white paper on 6G technologies for the wide-area cloud evolution. InterDigital and the University of Surrey announced a 6G research partnership to study joint communication and sensing and RIS technologies. India’s Department of Telecommunications reported the establishment of the Brahat 6G Alliance, consisting of 75 companies, with the goal of strengthening India’s 6G expertise and facilitating market access of Indian telecommunication products and services. In July, Northeastern University launched an Open Testing and Integration center to develop and test advanced O-RAN technologies. The GSM Association and the European Space Agency joined forces by signing an MoU to advance the integration of terrestrial and satellite networks in 5G and 6G. NTT and Fujitsu demonstrated a data rate of 30 Gbit/s over a 0.5 meter link distance at 300 GHz operating frequency using beamforming.
@@
In August, the Next G Alliance published 6G white papers on social/economic opportunities and spectrum considerations. The Next G Alliance also signed an MoU with the Indian Brahat 6G Alliance on 6G collaboration. In September, the NGMN Alliance released a 6G report from the perspective of network operators, titled ”6G Position Statement: An Operator View”. InterDigital and the Indian Institute of Technology, Kanpur agreed on a 6G cooperation with a special focus on extreme MIMO. LG Electronics and LG Uplus reported a record-breaking 6G spectrum test in an outdoor environment at LG Sciencepark in Seoul, sending data over 500 meters at THz frequencies. In October, Nokia opened a 6G laboratory at its Global R&\&&D center in Bangalore, India to promote 6G development, especially in the network sensing technology. InterDigital and the University Carlos III of Madrid agreed on a 6G research partnership, focusing on joint communication and sensing. Ericsson established a new 6G research team at its R&\&&D center in Chennai, India.
@@ I-D7 Activities in 2024
In January, ITU-R began the second phase of IMT-2030 development, i.e., defining its requirements and corresponding evaluation methodologies. 3GPP began the preparation of Release 19, continuing the development of 5G-Advanced toward 6G. NTT Docomo introduced their forthcoming AI-empowered 5G/6G technologies, such as human-augmented mobile platform to express immersive sensory perceptions in the metaverse environments.
@@
In February, SK Telecom and Intel reported that they have developed an AI-empowered ”Inline Service Mesh” technology for the 6G core network to reduce latency and increase service efficiency. China Mobile announced the launch of its 6G test satellite, claiming it to be the first of its kind in the world. Samsung and Princeton University announced their partnership in innovative 6G research through the Princeton’s NextG Initiative Corporate Affiliates Program, aiming to develop new technological innovations and foster cooperation between academia, industry, and regulators. Other collaborators in this program are Nokia Bell Labs, Ericsson, Vodafone, MediaTek, Intel, and Qualcomm Technologies. SK Telecom and Rohde &\&& Schwarz joined a group of industrial players collaborating in 6G spectrum trials at sub-THz frequencies. The existing partners include NTT Docomo, NTT, NEC Corporation, Fujitsu, Keysight Technologies, Nokia, and Ericsson. Nokia and the Indian Institute of Science announced their 6G collaboration, focusing on three 6G research areas: network architecture, radio technologies, and AI/ML-aided air interface. At the MWC 2024, Ericsson and Turkcell partnered in 6G research to promote technological advances in Turkie. Also at the MWC, Ericsson and Turk Telecom agreed on 6G cooperation by signing an MoU.
@@
In March, Nvidia announced that its 6G research cloud platform, based on a digital twin technology, is available, allowing simulations and testing in the realistic 6G environments. In April, Viavi Solutions reported its progress with Northeastern University in the development of a large-scale digital twin of a 6G network. The 6G Non-Terrestrial Networks project, funded by SNS JU, published a white paper on its vision of NTNs in 6G networks, highlighting the importance of non-terrestrial access and discussing the markets, coverage, and design of NTNs. The University of Glasgow reported that it has designed a digitally controlled metasurface antenna for the mmWave frequencies, paving the way toward advanced 6G antenna technologies. TDRA, the telecommunication regulator of the United Arab Emirates, unveiled its 6G roadmap, which promotes the nation’s leadership in the 6G development.

### v3

@@ I-D5 ACTIVITIES IN 2022
Many different types of activities occurred in January. The project TiC6G was launched by the University of Glasgow, with a group of academic and industrial partners, aiming to test prototype devices at THz frequencies using the new cutting-edge laboratory equipment. The Japanese operator Nippon Telegraph and Telephone (NTT) reported that it will build a 6G pilot network for the Osaka World Expo 2025. The Chinese Purple Mountain Laboratories, together with Fudan University and China Mobile, reported that it had obtained a data rate of 206.25 Gbit/s for a wireless THz transmission in a laboratory environment. The Next G Alliance published a white paper titled ”Green G: The Path toward Sustainable 6G[86]. The University of Oulu announced a partnership with Jio Estonia to study 6G technologies, particularly for non-terrestrial communications, holographic radio, and three-dimensional (3D) connected intelligence. Huawei published its 6G vision in a white paper entitled ”6G: The Next Horizon”, envisioning the transition from connected things to connected intelligence [87].
@@
In February, Keysight Technologies announced a partnership with Samsung to develop, test, and verify 6G technologies. The Next-Generation Mobile Networks (NGMN) Alliance published a white paper on 6G use cases, categorizing them into four main classes, i.e., enhanced human communications, enhanced machine communications, enabling services, and network evolution [88]. At the Mobile World Congress (MWC) 2022, LTE showcased its cutting-edge RIS technology, including four different prototypes, i.e., liquid crystal RIS, PIN diode RIS, transparent planar RIS, and transparent flexible RIS. At the same venue, the US technology company VMware reported its collaboration with German research institutes to accelerate 6G research, focusing on the fusion of cloud, networking, and AI technologies.
@@
In April, Kyocera, a Japanese technology firm, announced that it has developed a transmissive metasurface that can be used to extend coverage in 5G and 6G networks by redirecting radio signals to avoid obstacles. In May, an MoU was signed between the European 6G Smart Networks and Services Industry Association (6G-IA) and the Japanese Beyond 5G Promotion Consortium to accelerate the development of 6G networks and deepen the cooperation between Europe and Japan. Samsung released a new white paper, ”6G Spectrum: Expanding the Frontier[89], discussing spectrum policies and potential frequency bands for 6G. The Finnish leading academic and industrial organizations established a national 6G coalition, called 6G Finland, aiming to strengthen national 6G cooperation, form new international partnerships, and promote the Finnish 6G expertise on a global scale. LG demonstrated its 6G relevant full-duplex and power amplifier technologies at the IEEE ICC 2022. The Next G Alliance and the Japanese Beyond 5G Promotion Consortium agreed on 6G cooperation by signing an MoU in May.
@@
In June, the Next G Alliance published a white paper, titled ”6G Applications and Use Cases[86], focusing on four categories, i.e., networked robotics, XR, distributed communication and sensing, and personalized user experiences. Nokia, NTT, and NTT Docomo formed a 6G partnership to demonstrate an AI/ML-enhanced air interface and sub-THz communication at 140 GHz spectrum. The Singaporean Nanyang Technological University announced the use of Keysight test and measurement solutions to verify its on-chip THz electronic-photonic 6G device technologies. The Japanese companies NTT Docomo, Fujitsu, and NEC reported to perform joint 6G indoor and outdoor trials. 6G-IA and the Chinese IMT-2030 Promotion Group signed an MoU for 6G cooperation. The Open RAN (O-RAN) Alliance reported a launch of its Next Generation Research Group, focusing on the utilization of O-RAN technologies in 6G.
@@
In July, LG Uplus and Nokia signed an MoU for collaboration in 5G/6G standardization. Nokia announced its leadership in a German 6G light house project 6G-ANNA, aiming to promote 6G development and standardization in Germany. The Next G Alliance published a white paper, entitled ”6G Technologies[86], on the technologies required to meet the demands of its 6G vision. The UK government reported to grant 25 million pounds funding for the research and development work on 6G technologies. In August, the Next G Alliance and European 6G-IA signed an MoU for the agreement of 6G research collaboration, including joint workshops, webinars, and trials. While already supporting 6G research at the University of Texas at Austin, the Northeastern University, the University of Surrey, and Viavi Solutions announced a new funding program to advance worldwide 6G development in academia and industry. The Next G Alliance reported the establishment of a 6G research council, aiming to promote cooperation between academia and industry and align North American 6G strategy with the US and Canadian governments.
@@
In September, LG reported its final THz communication test at the Fraunhofer Heinrich-Hertz Institute in Berlin, with a successful information transfer over a 320 meter outdoor link at a frequency range between 155 and 175 GHz. The Finnish University of Oulu and Japanese National Institute of Communication and Technology (NICT) announced a cooperation in B5G/6G research. Rohde &\&& Schwarz and China Mobile reported their partnership in the development of ISAC solutions for 6G. The Singaporean Infocom Media Development Authority and the Singapore University of Technology and Design revealed their collaboration to establish a 6G Research and Development (R&\&&D) laboratory in Singapore, being the first in kind in Southeast Asia. Nokia informed its 5G/6G collaboration with Vodafone New Zealand. In October, Keysight Technologies and the Indian Institute of Technology, Madras signed an MoU, expanding their cooperation from 5G to 6G design. Deutsche Telekom reported its leadership in the German government-funded project 6G-TakeOff, aiming to develop an architectural framework for the integration of terrestrial and non-terrestrial networks in 6G. Ericsson and the University of Texas at Austin deepen their collaboration into 6G-empowered XR. VMware announced that it is launching the Next G-AI Research and Innovation Center in Montreal, Canada, gathering multidisciplinary expertise to advance the technological development of 5G and 6G systems. Samsung Electronics reported the founding of a new 6G research group in the UK, which is part of the global 6G development strategy. The German government-funded project 6G NeXt was kicked-off, with Deutsche Telekom in lead, and the goal of developing a flexible infrastructure platform to study and test future XR applications.
@@
In November, ITU-R WP 5D completed its work on the first report on IMT-2030, i.e., ”IMT-2030 Future Technology Trends” [19]. It was reported that the Indian Department of Telecommunications suggests opening up the frequency range from 95 GHz to 3 THz for experimenting and testing of 6G technologies and products. Nokia announced its leadership in the German government-supported KOMSENS-6G project, which focuses on the integration of sensing and communication. The Japanese Beyond 5G Promotion Group and Northeastern University agreed on a 6G collaboration by signing an MoU to foster the development of 6G networks. NTT Docomo and SK Telecom announced their partnership in the fields of smart-life and metaverse, searching for opportunities to jointly produce immersive XR content. Ericsson reported its ten-year 6G research initiative in the UK, focusing on the areas of AI, cognitive networks, and network security. Nokia announced the launch of a new 5G/6G R&\&&D center in Portugal. The UK government reported a 110 million pound grant for 5G/6G research and development. 28 million is dedicated to the collaboration of academia (Universities of York, Bristol, and Surrey) and industry (Nokia, Ericsson, and Samsung), while 80 million is invested in a new 5G/6G R&\&&D laboratory. All types of physical, hybrid, and virtual 6G events were held in 2022. Numerous IEEE special issues on 6G were released. A handful of 6G visions and surveys were also published [90, 91, 92, 93, 94, 95]. Moreover, lots of 6G white papers and books were written.
@@ I-D6 ACTIVITIES IN 2023
In January, the European Hexa-X-II flagship project (i.e., a continuation of Hexa-X) was launched [71]. As part of SNS JU Call 1, 35 other EU funded 5G/6G research and innovation projects were initiated, with a total funding of 250 million euros, in order to strengthen the European expertise in the development of future mobile networks [75]. 6G-IA and the European Telecommunications Standards Institute (ETSI) agreed on a close 5G/6G cooperation with each other by signing an MoU, with the goal of advancing 6G pre-standardization activities. The ATIS and O-RAN Alliance agreed on a partnership via an MoU to foster the development of O-RAN technologies. Rohde &\&& Schwarz announced its channel measurement campaign conducted in urban micro outdoor and indoor scenarios at frequencies of 158 GHz and 300 GHz, contributing to the ITU-R report titled ”Technical feasibility of IMT in bands above 100 GHz”. The Finnish University of Oulu and South Korean Electronics and Telecommunications Research Institute announced the launch of their new collaboration project, 6GBRIDGE-6GCORE, aiming to develop a service-centric 6G system architecture.
@@
In February, the University of Sheffield announced the opening of the new UK Research and Innovation National 6G Radio Systems Facility, with top-notch 6G R&\&&D capabilities and support from more than 40 companies and academic organizations. Nokia, NTT Docomo, and NTT reported two major milestones achieved on the evolution road toward 6G, including sub-THz communication (25 Gbit/s at 144 GHz) and the integration of AI into air interface (ML-based waveform), implemented as proof of concepts at Nokia Bell Labs in Germany. The Finnish 6G Flagship and Brazilian research institute Inatel announced their partnership in developing advanced 6G solutions for rural and remote areas. At the MWC 2023, Nokia demonstrated network sensing capabilities using its prototype radio equipment. Also at MWC, Bosch reported its collaboration with Nokia on 6G-based IoT solutions for industry 4.0 applications. The Taiwanese operator Chunghwa Telecom and Ericsson signed an MoU at MWC, joining forces to advance 5G and 6G standardization.
@@
In March, the Indian government unveiled a national 6G initiative, targeting to launch 6G networks in India by 2030. Anritsu and Danish Aalborg University announced their collaborative 6G project on channel sounding for joint communication and sensing at mmWave and sub-THz frequencies. In April, the Spanish operator Telefonica, NEC, Bluespecs, and IMDEA Research Institute reported the launch of the ENABLE-6G project, funded by the EU and Spanish government. The project aims to develop mechanisms to tackle the challenges 6G will encounter in terms of performance, energy efficiency, network sensing, and privacy/security. Keysight Technologies, the University of Surrey, and the National Physical Laboratory demonstrated over 100 Gbit/s data rates at a 300 GHz carrier frequency using their new sub-THz 6G testbed.
@@
In May, the Next G Alliance released a white paper on the 6G roadmap for vertical industries [86]. The French technology company Capgemini announced the opening of a new 6G research lab in India to enable testing and experimentation of novel 6G solutions. The Brazilian government reported the allocation of 36 million dollars to three research centers to advance the development of 5G, 6G, and O-RAN technologies. Rohde &\&& Schwarz and the French Institute of Electronics, Microelectronics and Nanotechnology announced their cooperation in THz communication research, with successful communication over 645 meters using a THz backhaul link operating at a frequency of 300 GHz.
@@
In June, the Next G Alliance published a white paper on 6G technologies for the wide-area cloud evolution [86]. InterDigital and the University of Surrey announced a 6G research partnership to study joint communication and sensing and RIS technologies. The Indian Department of Telecommunications reported the establishment of the Brahat 6G Alliance, consisting of 75 companies, with the goal of strengthening 6G expertise and facilitating market access of Indian telecommunication products and services. In July, Northeastern University launched an Open Testing and Integration center to develop and test advanced O-RAN technologies. The GSM Association and the European Space Agency joined forces by signing an MoU to advance the integration of terrestrial and satellite networks in 5G and 6G. NTT and Fujitsu demonstrated a data rate of 30 Gbit/s over a 0.5 meter link distance at 300 GHz operating frequency using beamforming.
@@
In August, the Next G Alliance published 6G white papers on social/economic opportunities and spectrum considerations [86]. The Next G Alliance also signed an MoU with the Indian Brahat 6G Alliance on 6G collaboration. In September, the NGMN Alliance released a 6G report from the perspective of network operators, titled ”6G Position Statement: An Operator View[88]. InterDigital and the Indian Institute of Technology, Kanpur agreed on a 6G cooperation with a special focus on extreme MIMO. LG Electronics and LG Uplus reported a record-breaking 6G spectrum test in an outdoor environment at LG Sciencepark in Seoul, sending data over 500 meters at THz frequencies. In October, Nokia opened a 6G laboratory at its Global R&\&&D center in Bangalore, India to promote 6G development, especially in the network sensing technology. InterDigital and the University Carlos III of Madrid agreed on a 6G research partnership, focusing on joint communication and sensing. Ericsson established a new 6G research team at its R&\&&D center in Chennai, India.
@@ I-D7 ACTIVITIES IN 2024
In January, ITU-R began the second phase of IMT-2030 development, i.e., defining its requirements and corresponding evaluation methodologies. 3GPP began the preparation of Release 19, continuing the development of 5G-Advanced toward 6G. SNS JU launched its Call 2 phase, including 27 6G research and innovation projects, with a total EU funding of 130 million euros [75]. NTT Docomo introduced their forthcoming AI-empowered 5G/6G technologies, such as human-augmented mobile platform to express immersive sensory perceptions in the metaverse environments.
@@
In February, SK Telecom and Intel reported that they have developed an AI-empowered ”Inline Service Mesh” technology for the 6G core network to reduce latency and increase service efficiency. China Mobile announced the launch of its 6G test satellite, claiming it to be the first of its kind in the world. Samsung and Princeton University announced their partnership in innovative 6G research through the Princeton NextG Initiative Corporate Affiliates Program, aiming to develop new technological innovations and foster cooperation between academia, industry, and regulators. Other collaborators in this program are Nokia Bell Labs, Ericsson, Vodafone, MediaTek, Intel, and Qualcomm Technologies. SK Telecom and Rohde &\&& Schwarz joined a group of industrial players collaborating in 6G spectrum trials at sub-THz frequencies. The existing partners include NTT Docomo, NTT, NEC Corporation, Fujitsu, Keysight Technologies, Nokia, and Ericsson. Nokia and the Indian Institute of Science announced their 6G collaboration, focusing on three 6G research areas: network architecture, radio technologies, and AI/ML-aided air interface. At the MWC 2024, Ericsson and Turkcell partnered in 6G research to promote technological advances in Turkie. Also at the MWC, Ericsson and Turk Telecom agreed on 6G cooperation by signing an MoU.
@@
In March, Nvidia announced that its 6G research cloud platform, based on a digital twin technology, is available, allowing simulations and testing in the realistic 6G environments. In April, Viavi Solutions reported its progress with Northeastern University in the development of a large-scale digital twin of a 6G network. The 6G Non-Terrestrial Networks project, funded by SNS JU, published a white paper on its vision of NTNs in 6G networks, highlighting the importance of non-terrestrial access and discussing the markets, coverage, and design of NTNs. The University of Glasgow reported that it has designed a digitally controlled metasurface antenna for the mmWave frequencies, paving the way toward advanced 6G antenna technologies. TDRA, the telecommunication regulator of the United Arab Emirates, unveiled its 6G roadmap, which promotes the leadership in the 6G development.